You are given that

  • Deaths are uniformly distributed over each year of age.
  • LaTeX: Z Z is the p.v.r.v. of a LaTeX: m m -thly LaTeX: n n -year endowment contract.
  • The benefit of the contract is 1.
  • LaTeX: m=3 m = 3 and n-2 n = 2
  • і 0.06 i = 0.06
  • 1 0 q x = 0.1 and +1 =0.15 q x + 1 = 0.15

Calculate E(Z) E ( Z ) .





n-2
і 0.06
1 0
+1 =0.15
E(Z)
Answer #1

the 2 09434 = dog434 + 0 120 150 Crm Efz产ー.vn 낙-6 E(Z)= I-(09934 ) O 0230 3.9153 E(T) =